# Recursive subhomogeneity of orbit-breaking subalgebras of $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles

### v1

@@ 6. Endomorphism bundles
In this section we construct endomorphism bundles vector bundles constructed from line bundles. These will be the buildings blocks in the recursive subhomogeneous decomposition of the orbit-breaking algebra. We discuss the local trivialization of these bundles in detail.
First we recall that for a vector bundle $\mathscr{W}$ of rank $n$ over $X$, the endomorphism bundle $End(\mathscr{W})$ is a vector bundle over $X$ whose fibre at $x$ consists of all linear maps on the fibre vector space $\mathscr{W}_{x}$ of $\mathscr{W}$, namely $End(\mathscr{W})_{x}=End(\mathscr{W}_{x})$ for every $x$ (see [20, Example 6.8]).
Let $\mathcal{U}$ be an open cover of $X$ and let $\{(U,h_{U}:U\times\mathbb{C}^{n}\to\mathscr{W}|_{U})\}_{U\in\mathcal{U}}$ an atlas of $\mathscr{W}$ with transition functions $\{g_{U,V}\}_{U,V\in\mathcal{U}}$. Then it induces an atlas $\{(U,H_{U}:U\times M_{n}(\mathbb{C})\to End(\mathscr{W})|_{U})\}$ of $End(\mathscr{W})$ as follows:
@@
for $M\in M_{n}(\mathbb{C})$, which is represented in the following diagram with $h_{U,x}:=h_{U}(x,\cdot)$.
@@
is well defined and it satisfies
@@
$=\left(x,\mathrm{Ad}(g_{V,U}(x))M\right).$
@@ 6.0.1. Endomorphism bundles $\mathscr{M}^{(n)}$
Let $\mathscr{V}$ be a line bundle over $X$. Let $\mathscr{D}^{(m,n)}$ and $\mathscr{D}^{(n)}$ be vector bundles over $X$ given by
@@
Since $\mathscr{D}^{(n)}$ is of rank $n$, the bundle $\mathscr{M}^{(n)}$ has fiber $\mathscr{M}^{(n)}_{x}=End(\mathscr{D}^{(n)}_{x})$ isomorphic to $M_{n}(\mathbb{C})$.
@@ 6.0.2. Local trivialization of $\mathscr{M}^{(n)}$
We fix a finite open cover $\mathcal{U}_{0}$ of $X$ such that $\mathscr{V}$ is trivial over $\alpha^{i}(U)$ with a chart map
$h_{\alpha^{i}(U)}:\alpha^{i}(U)\times\mathbb{C}\to\mathscr{V}|_{\alpha^{i}(U)}$
for all $U\in\mathcal{U}_{0}$ and $0\leq i\leq 2n$. Then the atlas
$\{(\alpha^{i}(U),h_{\alpha^{i}(U)})\mid U\in\mathcal{U}_{0},\,i=0,\dots,2n\}$
gives us transition functions $g_{\alpha^{i}(U),\alpha^{j}(V)}:\alpha^{i}(U)\cap\alpha^{j}(V)\to U(1)$ such that
$h_{\alpha^{i}(U)}(x,\lambda)=h_{\alpha^{j}(V)}(x,g_{\alpha^{j}(V),\alpha^{i}(U)}(x)\lambda),\ \lambda\in\mathbb{C},$ (10)
for all $U\in\mathcal{U}_{0}$, $0\leq i\leq 2n$. Let
@@
Then for the line bundle $\mathscr{V}^{(l)}$, $10\leq l\leq n-$ 1, the following maps
$(x,\lambda)\mapsto\lambda\,h_{\alpha^{l-1}(U)}(\alpha^{l-1}(x),1)\otimes\cdots\otimes h_{\alpha(U)}(\alpha(x),1)\otimes h_{U}(x,1)$
from $U\times\mathbb{C}$ onto $\mathscr{V}^{(l)}|_{U}$, $U\in\mathcal{U}$, form an atlas. For convenience, we set
$v^{(l)}_{U}(x):=h_{\alpha^{l-1}(U)}(\alpha^{l-1}(x),1)\otimes\cdots\otimes h_{\alpha(U)}(\alpha(x),1)\otimes h_{U}(x,1),\ x\in U.$
Then $v_{U}^{(l)}:U\to\mathscr{V}^{(l)}|_{U}$ is a nonzero local section and thus each of the vectors $v^{(l)}_{U}(x)$, $U\in\mathcal{U}$ with $x\in U$, serves as a basis element of the one dimensional space $\mathscr{V}^{(l)}_{x}$. By (10) $v^{(l)}_{U}(x)$ and $v^{(l)}_{V}(x)$ satisfy the relation
$v^{(l)}_{U}(x)=\Pi_{i=0}^{l-1}g_{\alpha^{j}(V),\alpha^{j}(U)}(\alpha^{j}(x))\,v^{(l)}_{V}(x)$
whenever $x\in U\cap V$ for $U,V\in\mathcal{U}$. In turn, one has a basis of the $n$-dimensional space $\mathscr{D}^{(n)}_{x}=\mathscr{V}^{(0)}_{x}\oplus\mathscr{V}^{(1)}_{x}\oplus\cdots\oplus\mathscr{V}^{(n-1)}_{x}$ at each $x\in U$ given by
$e_{U,0}(x)$ $:=(v_{U}^{(0)}(x),0,0,\ldots,0),$
$e_{U,1}(x)$ $:=(0,v_{U}^{(1)}(x),0,\dots,0),$
@@
$e_{U,n-1}(x)$ $:=(0,0,\dots,0,v_{U}^{(n-1)}(x))$
with convention $v_{U}^{(0)}(x):=1$. Then
$(\lambda_{0},\dots,\lambda_{n-1})\mapsto\lambda_{0}e_{U,0}(x)+\cdots+\lambda_{n-1}e_{U,n-1}(x):\mathbb{C}^{n}\to\mathscr{D}_{x}^{(n)},\ x\in U,$
gives us a chart map of $\mathscr{D}^{(n)}_{U}$ (by continuity of $e_{U,i}$’s over $U$) from which we obtain a chart map
$H_{U}^{(n)}:U\times M_{n}(\mathbb{C})\to\mathscr{M}_{U}^{(n)}:=\mathscr{M}^{(n)}|_{U}$
of the endomorphism bundle $\mathscr{M}^{(n)}=End(\mathscr{D}^{(n)})$ over $U\in\mathcal{U}$. Hence for a linear map $\varsigma(x)\in\mathscr{M}^{(n)}_{x}$, $x\in U$, there exists a unique matrix $\widetilde{\varsigma}_{U}(x)\in M_{n}(\mathbb{C})$ such that $H_{U}^{(n)}(x,\widetilde{\varsigma}_{U}(x))=\varsigma(x)$. Namely, if $\widetilde{\varsigma}_{U}(x)$ is written as
$\widetilde{\varsigma}_{U}(x)=\footnotesize{\begin{bmatrix}a_{00}(x)&a_{01}(x)&a_{02}(x)&\cdots&a_{0(n-1)}(x)\\ a_{10}(x)&a_{11}(x)&a_{12}(x)&\cdots&a_{1(n-1)}(x)\\ a_{20}(x)&a_{21}(x)&a_{22}(x)&\cdots&a_{2(n-1)}(x)\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{(n-1)0}(x)&a_{(n-1)1}(x)&a_{(n-1)2}(x)&\cdots&a_{(n-1)(n-1)}(x)\end{bmatrix}},$
then it means that for each $i=0,\dots,n-1$,
$\varsigma(x)e_{U,i}(x)=a_{0i}(x)e_{U,0}(x)+a_{1i}(x)e_{U,1}(x)+\cdots+a_{(n-1)i}(x)e_{U,n-1}(x).$
When $x$ is fixed and clear in context, by abusing notation, we write simply $\widetilde{\varsigma}_{U}(x)=(H_{U}^{(n)})^{-1}(\varsigma(x))$.
Now we consider what happens to the matrix representation of a linear map in $\mathscr{M}_{x}^{(n)}$ when we make a coordinate change in the vector space $\mathscr{D}^{(n)}_{x}$ for $x$ in an overlap $U\cap V$. For this, first set
$g_{VU}^{(l)}(x):=\prod_{i=0}^{l-1}{g_{\alpha^{i}(V),\alpha^{i}(U)}}(\alpha^{i}(x))$
and consider the following diagonal unitary matrix
$u_{VU}(x):=diag(1,g_{VU}^{(1)}(x),g_{VU}^{(2)}(x),\ldots,g_{VU}^{(n-1)}(x)).$ (11)
@@
$\widetilde{\varsigma}_{V}(x)=u_{VU}(x)\widetilde{\varsigma}_{U}(x)u_{VU}^{-1}(x).$ (12)
@@
$H_{U}^{(n)}(x,\widetilde{\varsigma}_{U}(x))=H_{V}^{(n)}(x,\mathrm{Ad}(u_{V,U}(x))\widetilde{\varsigma}_{U}(x)),$
and $x\mapsto\mathrm{Ad}(u_{V,U}(x)):X\to Aut(M_{n}(\mathbb{C}))$ are the transition functions for the atlas $\{(U,H_{U}^{(n)})\}_{U\in\mathcal{U}}$ of $\mathscr{M}^{(n)}$, where $\mathrm{Ad}(u)$ is an automorphism such that $\mathrm{Ad}(u)M=uMu^{*}$ for $M\in M_{n}(\mathbb{C})$.
Definition 6.1.
We call the atlas $\{(U,H_{U}^{(n)})\}_{U\in\mathcal{U}}$ the local trivialization induced from the atlas $\{h_{\alpha^{j}(U)}:\alpha^{j}(U)\times\mathbb{C}\to\mathscr{V}|_{\alpha^{j}(U)}\mid U\in\mathcal{U},\ j=0,\dots,n\}$.
If $\widetilde{\varsigma}_{U}(x)$ is written as above then with $(x)$ omitted,
$\widetilde{\varsigma}_{V}$ $=u_{VU}\widetilde{\varsigma}_{U}{u_{VU}}^{-1}$
$=\footnotesize{\begin{bmatrix}a_{00}&a_{01}{g_{VU}^{(1,0)}}^{-1}&a_{02}{g_{VU}^{(2,0)}}^{-1}&\cdots&a_{0(n-1)}{g_{VU}^{(n-1,0)}}^{-1}\\ g_{VU}^{(1,0)}a_{10}&a_{11}&a_{12}{g_{VU}^{(2,1)}}^{-1}&\cdots&a_{1(n-1)}{g_{VU}^{(n-1,1)}}^{-1}\\ g_{VU}^{(2,0)}a_{20}&{g_{VU}^{(2,1)}}a_{21}&a_{22}&\cdots&a_{2(n-1)}{g_{VU}^{(n-1,2)}}^{-1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ g_{VU}^{(n-1,0)}a_{(n-1)0}&{g_{VU}^{(n-1,1)}}a_{(n-1)1}&{g_{VU}^{(n-1,2)}}a_{(n-1)2}&\cdots&a_{(n-1)(n-1)}\end{bmatrix}},$
where $g_{VU}^{(k,0)}(x):=g_{VU}^{(k)}(x)$ and $g_{VU}^{(k,j)}(x):=\prod_{i=j}^{k-1}{g_{\alpha^{i}(V)\alpha^{i}(U)}}(\alpha^{i}(x))$. Note that the matrices $\widetilde{\varsigma}_{U}(x)$ and $\widetilde{\varsigma}_{V}(x)$ have the same diagonal elements, that is, the diagonal is invariant under the coordinate change from $\{e_{U,i}(x)\}$ to $\{e_{V,i}(x)\}$ for $x\in U\cap V$. Also if one of the matrices is $m$th subdiagonal, then so is the other for $11\leq m\leq n-$ 1.
By a matrix representation of a linear map $\varsigma(x)$ in $\mathscr{M}^{(n)}_{x}$, we always mean a matrix $\widetilde{\varsigma}_{U}(x)=(H_{U}^{(n)})^{-1}(\varsigma(x))$ for some $U\in\mathcal{U}$ with $x\in U$ if not mentioned otherwise. We have the following remark from the above observations.
@@
Let $\varsigma\in\Gamma(\mathscr{M}^{(n)})$ and let $\widetilde{\varsigma}_{U}(x)=(H_{U}^{(n)})^{-1}(\varsigma(x))$ for $U\in\mathcal{U}$ and $x\in U$. Then we have the following:
@@
    The main diagonal is invariant: $\mathrm{diag}(\widetilde{\varsigma}_{U}(x))=\mathrm{diag}(\widetilde{\varsigma}_{V}(x))$ whenever $x\in U\cap V$.
@@
    If $\widetilde{\varsigma}_{U}(x)$ is $m$th subdiagonal, then so is $\widetilde{\varsigma}_{V}(x)$ whenever $x\in U\cap V$.
@@
    If $\varsigma(x)$ is represented as an $m$th (lower) subdiagonal matrix $\widetilde{\varsigma}_{U}(x)=(H_{U}^{(n)})^{-1}(\varsigma(x))$ on $U$ for $11\leq m\leq n-$ 1, then $\widetilde{\varsigma}_{U}(x)$ is of the form
@@
    where each $s_{i,j}$ is a continuous function on $U$. In this case $\varsigma(x)$ maps each element $e_{U,i}(x)$ in basis of $\mathscr{D}^{(n)}_{x}$ to
    $\varsigma(x)e_{U,i}(x)=\begin{cases}s_{m+i,i}(x)e_{U,m+i}(x),&\text{for $0\leq i<n-m$},\\ 0,&\text{for $n-m\leq i\leq n-1$}.\end{cases}$ (13)
    We recall the definition of the basis element $e_{U,i}(x)$ and see that $\varsigma(x)$ maps each $\mathscr{V}^{(i)}_{x}$ to $\mathscr{V}^{(i+m)}_{x}$ for $0\leq i<n-m$, and furthermore, the linear map

### v2

@@ 6. Endomorphism bundles
In this section we construct endomorphism bundles of vector bundles constructed from line bundles. The section algebras of these endomorphism bundles will be the building blocks in the recursive subhomogeneous decomposition of the orbit-breaking algebra. We discuss the local trivialization of these bundles in detail.
First we recall that for a vector bundle $\mathscr{W}$ of rank $n$ over $X$, the endomorphism bundle $End(\mathscr{W})$ is a vector bundle over $X$ whose fibre at $x$ consists of all linear maps on the fibre vector space $\mathscr{W}_{x}$ of $\mathscr{W}$, namely $End(\mathscr{W})_{x}=End(\mathscr{W}_{x})$ for every $x$ (see [22, 6.8 Example]).
Let $\mathcal{U}$ be an open cover of $X$ and let $\{(U,h_{U}:U\times\mathbb{C}^{n}\to\mathscr{W}|_{U})\}_{U\in\mathcal{U}}$ be an atlas of $\mathscr{W}$ with transition functions $\{g_{U,V}\}_{U,V\in\mathcal{U}}$. Then $\mathcal{U}$ induces an atlas $\{(U,H_{U}:U\times M_{n}(\mathbb{C})\to End(\mathscr{W})|_{U})\}$ of $End(\mathscr{W})$ where the $H_{U}$ are defined by
@@
for $M\in M_{n}(\mathbb{C})$. This is represented in the following diagram with $h_{U,x}:=h_{U}(x,\cdot)$:
@@
is well defined and satisfies
@@
$=\left(x,\mathrm{Ad}(g_{V,U}(x))M\right),$
where $\mathrm{Ad}(u)$ for a unitary $u\in M_{n}(\mathbb{C})$ denotes the automorphism of $M_{n}(\mathbb{C})$ such that $\mathrm{Ad}(u)M=uMu^{*}$.
@@ 6.0.1. Endomorphism bundles $\mathscr{M}^{(n)}$
Let $\mathscr{V}$ be a line bundle over $X$ and $\alpha:X\to X$ a homeomorphism. With
$\mathscr{V}^{(n)}:=(\alpha^{n-1})^{*}\mathscr{V}\otimes(\alpha^{n-2})^{*}\mathscr{V}\otimes\cdots\otimes\alpha^{*}\mathcal{\mathscr{V}}\otimes\mathscr{V}$
as in (3), let $\mathscr{D}^{(m,n)}$ and $\mathscr{D}^{(n)}$ be vector bundles over $X$ given by
@@
Since $\mathscr{D}^{(n)}$ is of rank $n$, the bundle $\mathscr{M}^{(n)}$ has fibre $\mathscr{M}^{(n)}_{x}=End(\mathscr{D}^{(n)}_{x})$ isomorphic to $M_{n}(\mathbb{C})$.
@@ 6.0.2. Local trivialization of $\mathscr{M}^{(n)}$
We fix a finite open cover $\mathcal{U}=\{U_{j}\}_{j}$ of $X$ such that $\mathscr{V}$ is trivial over $U_{j}$ with a chart map
$h_{U_{j}}:U_{j}\times\mathbb{C}\to\mathscr{V}|_{U_{j}}$
for each $U_{j}\in\mathcal{U}$. From the atlas $\{(U_{j},h_{U_{j}})\mid U_{j}\in\mathcal{U}\}$ we obtain the transition functions $g_{U_{i},U_{j}}:U_{i}\cap U_{j}\to U(1)$ such that
$h_{U_{i}}(x,\lambda)=h_{U_{j}}(x,g_{U_{j},U_{i}}(x)\lambda)$ (10)
for all $x\in U_{i}\cap U_{j}$ and $\lambda\in\mathbb{C}$. Fix $n\in\mathbb{N}$ and let
$\mathcal{U}^{(n)}:=\{(U_{0},U_{1},\dots,U_{n-1})\mid U_{j}\in\mathcal{U}\text{and }\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})\neq\emptyset\}.$
Note that if $n=1$, then $\mathcal{U}^{(1)}=\mathcal{U}$, and that
$x\in\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})\ \Leftrightarrow\,\alpha^{j}(x)\in U_{j}\ \text{ for all }j=1,\dots,n-1.$
For convenience, we say that $x$ is contained in $\mathbf{U}=(U_{0},\dots,U_{n-1})\in\mathcal{U}^{(n)}$ if $x\in\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})$. In this case, we simply write $x\in\mathbf{U}$:
$x\in\mathbf{U}=(U_{0},\dots,U_{n-1})$ $\ \Leftrightarrow\ x\in\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})$
$\ \Leftrightarrow\ \alpha^{j}(x)\in U_{j}\ \text{for all }j=0,\dots,n-1.$
For $\mathbf{U}\in\mathcal{U}^{(n)}$ and $1\leq l\leq n$, the map
$v^{(l)}_{\mathbf{U}}:\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})\to\mathscr{V}^{(l)}|_{\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})}$
given by
$v^{(l)}_{\mathbf{U}}(x):=h_{U_{l-1}}(\alpha^{l-1}(x),1)\otimes\cdots\otimes h_{U_{1}}(\alpha(x),1)\otimes h_{U_{0}}(x,1)$
is a non-zero continuous local section of the line bundle $\mathscr{V}^{(l)}$ over the open set $\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})$. The set of all open sets $\{\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})\mid\mathbf{U}=(U_{0},\dots,U_{n-1})\in\mathcal{U}^{(n)}\}$ is a finite open cover of $X$ and forms an atlas of $\mathscr{V}^{(l)}$ together with the following maps
$(x,\lambda)\mapsto\lambda v^{(l)}_{\mathbf{U}}(x):\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})\times\mathbb{C}\to\mathscr{V}^{(l)}|_{\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})}.$
For later use, we set up some notation: If $x\in\mathbf{U}\in\mathcal{U}^{(n)}$, then $\alpha^{i}(x)\in(U_{i},\dots,U_{n-1})\in\mathcal{U}^{(n-i)}$ for $i=0,\dots,n-1$. Reflecting this point we will use the following notation:
$\alpha^{i}(\mathbf{U}):=(U_{i},\dots,U_{n-1})\in\mathcal{U}^{(n-i)}.$
Also for $11\leq m\leq n-$ 1 and $10\leq i\leq n-$ 1 with $1\leq m+i\leq n$, we write
$v^{(m)}_{\alpha^{i}(\mathbf{U})}(\alpha^{i}(x)):=h_{U_{m+i-1}}(\alpha^{m+i-1}(x),1)\otimes\cdots\otimes h_{U_{i}}(\alpha^{i}(x),1).$ (11)
Since $v_{\mathbf{U}}^{(l)}$ is a non-zero local section of $\mathscr{V}^{(l)}$, each unit vector $v_{\mathbf{U}}^{(l)}(x)$ serves as a basis element of the one dimensional space $\mathscr{V}^{(l)}_{x}$ for all $x\in\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})$. In turn, one has a basis of the $n$-dimensional space $\mathscr{D}^{(n)}_{x}=\mathscr{V}^{(0)}_{x}\oplus\mathscr{V}^{(1)}_{x}\oplus\cdots\oplus\mathscr{V}^{(n-1)}_{x}$ at $x\in\mathbf{U}\in\mathcal{U}^{(n)}$ given by
@@
[…]
with the convention that $v_{\mathbf{U}}^{(0)}(x):=1$. Then the map
$(x,(\lambda_{0},\dots,\lambda_{n-1}))\mapsto\lambda_{0}\,e_{\mathbf{U},0}(x)+\cdots+\lambda_{n-1}e_{\mathbf{U},n-1}(x)$
from $\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})\times\mathbb{C}^{n}$ onto $\mathscr{D}^{(n)}|_{\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})}$ is a chart map of $\mathscr{D}^{(n)}$ (by continuity of $e_{\mathbf{U},i}$’s over $\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})$) from which we obtain a chart map
$H_{\mathbf{U}}^{(n)}:\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})\times M_{n}(\mathbb{C})\to\mathscr{M}_{\mathbf{U}}^{(n)}:=\mathscr{M}^{(n)}|_{\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})}$
of the endomorphism bundle $\mathscr{M}^{(n)}=End(\mathscr{D}^{(n)})$ over $\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})$. Hence, for a linear map $\varsigma(x)\in\mathscr{M}^{(n)}_{x}$, $x\in\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})$, there exists a unique matrix $\widetilde{\varsigma}_{\mathbf{U}}(x)\in M_{n}(\mathbb{C})$ such that $H_{\mathbf{U}}^{(n)}(x,\widetilde{\varsigma}_{\mathbf{U}}(x))=\varsigma(x)$. Namely, if $\widetilde{\varsigma}_{\mathbf{U}}(x)$ is written as
$\widetilde{\varsigma}_{\mathbf{U}}(x)=\scriptstyle{\begin{bmatrix}a_{00}(x)&a_{01}(x)&a_{02}(x)&\cdots&a_{0(n-1)}(x)\\ a_{10}(x)&a_{11}(x)&a_{12}(x)&\cdots&a_{1(n-1)}(x)\\ a_{20}(x)&a_{21}(x)&a_{22}(x)&\cdots&a_{2(n-1)}(x)\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{(n-1)0}(x)&a_{(n-1)1}(x)&a_{(n-1)2}(x)&\cdots&a_{(n-1)(n-1)}(x)\end{bmatrix}},$
then for each $j=0,\dots,n-1$,
$\varsigma(x)e_{\mathbf{U},j}(x)\\ =a_{0j}(x)e_{\mathbf{U},0}(x)+a_{1j}(x)e_{\mathbf{U},1}(x)+\cdots+a_{(n-1)j}(x)e_{\mathbf{U},n-1}(x).$
When $x$ is fixed and clear in context, by abuse of notation, we simply write $\widetilde{\varsigma}_{\mathbf{U}}(x)=(H_{\mathbf{U}}^{(n)})^{-1}(\varsigma(x))$.
Definition 6.1.
We call the atlas $\{\big{(}\cap_{j=0}^{n-1}\alpha^{-j}(U_{j}),H_{\mathbf{U}}^{(n)}\big{)}\}_{\mathbf{U}\in\mathcal{U}^{(n)}}$ the local trivialization of $\mathscr{M}^{(n)}$ induced from the atlas $\{(U_{j},h_{U_{j}})\}_{U_{j}\in\mathcal{U}}$ of $\mathscr{V}$.
Now we consider what happens to the matrix representation of a linear map in $\mathscr{M}_{x}^{(n)}$ when we make a coordinate change in the vector space $\mathscr{D}^{(n)}_{x}$. For this, let $\mathbf{U},\mathbf{V}\in\mathcal{U}^{(n)}$, $x\in\mathbf{U}\cap\mathbf{V}$ and $0\leq l\leq n$. Then we have from (10) that
$\ v^{(l)}_{\mathbf{U}}(x)$
$=$ $\ h_{U_{l-1}}(\alpha^{l-1}(x),1)\otimes\cdots\otimes h_{U_{1}}(\alpha(x),1)\otimes h_{U_{0}}(x,1)$
$=$ $\ \Pi_{j=0}^{l-1}g_{V_{j},U_{j}}(\alpha^{j}(x))\cdot h_{V_{l-1}}(\alpha^{l-1}(x),1)\otimes\cdots\otimes h_{V_{1}}(\alpha(x),1)\otimes h_{V_{0}}(x,1)$
$=$ $\ g^{(l)}_{\mathbf{V},\mathbf{U}}(x)\cdot v^{(l)}_{\mathbf{V}}(x),$
where $g^{(l)}_{\mathbf{V},\mathbf{U}}(x):=\Pi_{j=0}^{l-1}\,g_{V_{j},U_{j}}(\alpha^{j}(x)).$ With the following diagonal unitary matrix
$u_{\mathbf{VU}}(x):=\mathrm{diag}(1,g_{\mathbf{V},\mathbf{U}}^{(1)}(x),g_{\mathbf{V},\mathbf{U}}^{(2)}(x),\ldots,g_{\mathbf{V},\mathbf{U}}^{(n-1)}(x)),$ (12)
@@
$\widetilde{\varsigma}_{\mathbf{V}}(x)=u_{\mathbf{VU}}(x)\widetilde{\varsigma}_{\mathbf{U}}(x)u_{\mathbf{VU}}^{-1}(x).$ (13)
@@
$H_{\mathbf{U}}^{(n)}(x,\widetilde{\varsigma}_{\mathbf{U}}(x))=H_{\mathbf{V}}^{(n)}(x,\mathrm{Ad}(u_{\mathbf{VU}}(x))\widetilde{\varsigma}_{\mathbf{U}}(x)),$
and $x\mapsto\mathrm{Ad}(u_{\mathbf{VU}}(x)):\cap_{j=0}^{n-1}\alpha^{-j}(U_{j}\cap V_{j})\to\mathrm{Aut}(M_{n}(\mathbb{C}))$ are the transition functions for the atlas $\{\big{(}\cap_{j=0}^{n-1}\alpha^{-j}(U_{j}),H_{\mathbf{U}}^{(n)}\big{)}\}_{\mathbf{U}\in\mathcal{U}^{(n)}}$ of $\mathscr{M}^{(n)}$.
If $\widetilde{\varsigma}_{\mathbf{U}}(x)$ is written as above, then with $(x)$ omitted,
$\ \widetilde{\varsigma}_{\mathbf{V}}=\ u_{\mathbf{VU}}\widetilde{\varsigma}_{\mathbf{U}}{u_{\mathbf{VU}}}^{-1}$
$=$ $\ \scriptstyle{\begin{bmatrix}a_{00}&a_{01}{g_{\mathbf{V},\mathbf{U}}^{(1)}}^{-1}&\cdots&a_{0(n-1)}{g_{\mathbf{V},\mathbf{U}}^{(n-1)}}^{-1}\\ g_{\mathbf{V,U}}^{(1)}a_{10}&a_{11}&\cdots&g_{\mathbf{V,U}}^{(1)}a_{1(n-1)}{g_{\mathbf{V},\mathbf{U}}^{(n-1)}}^{-1}\\ \vdots&\vdots&\ddots&\vdots\\ g_{\mathbf{V},\mathbf{U}}^{(n-1)}a_{(n-1)0}&{g_{\mathbf{V},\mathbf{U}}^{(n-1)}}a_{(n-1)2}{g_{\mathbf{V},\mathbf{U}}^{(2)}}^{-1}&\cdots&a_{(n-1)(n-1)}\end{bmatrix}}.$
Note that the matrices $\widetilde{\varsigma}_{\mathbf{U}}(x)$ and $\widetilde{\varsigma}_{\mathbf{V}}(x)$ have the same diagonal elements, that is, the diagonal is invariant under the coordinate change from $\{e_{\mathbf{U},i}(x)\}$ to $\{e_{\mathbf{V},i}(x)\}$ for $x\in\mathbf{U}\cap\mathbf{V}$. Also if one of the matrices is $m$th subdiagonal, then so is the other for $11\leq m\leq n-$ 1.
By a matrix representation of a linear map $\varsigma(x)$ in $\mathscr{M}^{(n)}_{x}$, we always mean a matrix $\widetilde{\varsigma}_{U}(x)=(H_{\mathbf{U}}^{(n)})^{-1}(\varsigma(x))$ for some $\mathbf{U}=(U_{0},\dots,U_{n-1})\in\mathcal{U}^{(n)}$ with $x\in\mathbf{U}$ if not mentioned otherwise. We have the following remark from the above observations.
@@
Let $\varsigma\in\Gamma(\mathscr{M}^{(n)})$ and let $\widetilde{\varsigma}_{\mathbf{U}}(x)=(H_{\mathbf{U}}^{(n)})^{-1}(\varsigma(x))\in M_{n}(\mathbb{C})$ be a matrix representation of $\varsigma(x)$ for $x$ contained in $\mathbf{U}\in\mathcal{U}^{(n)}$. Then we have the following:
@@
    The main diagonal is invariant: $\mathrm{diag}(\widetilde{\varsigma}_{\mathbf{U}}(x))=\mathrm{diag}(\widetilde{\varsigma}_{\mathbf{V}}(x))$ whenever $x\in\mathbf{U}\cap\mathbf{V}$ for $\mathbf{U,V}\in\mathcal{U}^{(n)}$.
@@
    If $\widetilde{\varsigma}_{\mathbf{U}}(x)$ is $m$th subdiagonal, then so is $\widetilde{\varsigma}_{\mathbf{V}}(x)$ whenever $x\in\mathbf{U}\cap\mathbf{V}$.
@@
    If $\varsigma(x)$ is represented as an $m$th (lower) subdiagonal matrix $\widetilde{\varsigma}_{\mathbf{U}}(x)=(H_{\mathbf{U}}^{(n)})^{-1}(\varsigma(x))$ on $\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})$ for $11\leq m\leq n-$ 1, then $\widetilde{\varsigma}_{\mathbf{U}}(x)$ is of the form
@@
    where each $s_{i,j}$ is a continuous function on $\cap_{j=0}^{n-1}\alpha^{-j}(U_{j})$. In this case $\varsigma(x)$ maps each element $e_{\mathbf{U},i}(x)$ in basis of $\mathscr{D}^{(n)}_{x}$ to
    $\varsigma(x)e_{\mathbf{U},i}(x)=\begin{cases}s_{m+i,i}(x)e_{\mathbf{U},m+i}(x),&\text{for $0\leq i<n-m$},\\ 0,&\text{for $n-m\leq i\leq n-1$}.\end{cases}$ (14)
    We recall the definition of the basis element $e_{\mathbf{U},i}(x)$ and see that $\varsigma(x)$ maps each $\mathscr{V}^{(i)}_{x}$ to $\mathscr{V}^{(i+m)}_{x}$ for $0\leq i<n-m$, and furthermore, the linear map